10) If a eukaryotic virus is a 60 nm diameter sphere, about how many will fit into a eukaryotic cell above?
Answer #1

Number of virus in eukaryotic cell = volume of eukaryotic cell / volume of virus

Diameter of eukaryotic cell = 100 um = 100 × 1000 = 10^5 nm (maximum diameter)

Volume of eukaryotic cell = 4/3 π r^3 = 4/3 (3.14) (10^5)^3 = 4.186 × 10^15 nm^3

Volume of virus = 4/3 π r^3 = 4/3 (3.14) (60)^3 = 9.04 × 10^5 nm^3

Number of viruses 4.186 × 10^15 / 9.04 × 10^5 = 4.63 × 10^9 cells

Please rate.